# Relation Extraction with Instance-Adapted Predicate Descriptions

### v1

@@ Methods
During our implementation, though contrastive training (from Equation (Methods)) was effective in pushing the input and positive predicate representations closer, we observed it does not learn robust relation representations of the input (the s). To address this, we propose to add a new linear layer and simultaneously optimize the cross-entropy loss
@@
Assessing different description texts: Although we used predicate definitions created by RE dataset creators, they are the nevertheless a single symbolic and discrete form. We wondered about how the scores would change if we rephrased them by prompting GPT-4 to “rewrite” them without any special instructions. We also tested a simple baseline that literally has the predicate name and entity spans fill the slots in this template: @predicate@: @subject@, @object@.
@@
Next, consider this input from the ChemProt dataset: “Down-regulation of prostate-specific antigen (PSA) expression, an AR-target gene, by estramustine and bicalutamide was accompanied by the blockade of the mutated androgen receptor.” The model was able to identify the INDIRECT-DOWNREGULATOR relation between drugs estramustine and bicalutamide (subjects) and the protein PSA (object). But it failed to spot the same relation of those drugs with the object protein AR, which appears to be implicitly stated. Considering ChemProt extraction sometimes invovles the full abstract, other sentences surrounding this may offer indirect clues about the relation. However, upon examining the full abstract, we are not able to see stronger evidence than what is already present in the sentence shown here.
Another common error pattern is due to the model’s lack of grasp of deep domain knowledge, particularly in biomedical datasets when information is densely packed. For example, consider this pithy ChemProt input: “In vivo, agonist actions of yohimbine at 5-HT(1A) sites are revealed by WAY 100,635-reversible induction of hypothermia in the rat.” Here the gold relation is the ANTAGONIST link between subject WAY 100,635 and object 5-HT(1A). The agonist actions of the chemical yohimbine on 5-HT(1A) result in hypothermia and the fact that this hypothermia can be reversed by WAY 100,635 indicates that it is playing an antagonist role for 5-HT(1A). This a complex expression involving the intermediate entity yohimbine and an unusual looking phrase WAY 100,635-reversible that densely packs meaning that typically needs a new sentence to convey explicitly. This may simply be a case of a highly complex example needing multi-hop reasoning that needed to be carried out with a compact input text.

### v2

@@ Methods
During our implementation, though contrastive training (from Equation (3)) was effective in pushing the input and positive predicate representations closer, we observed it does not learn robust relation representations of the input (the s). To address this, we propose to add a new linear layer and simultaneously optimize the cross-entropy loss
@@
Assessing different descriptions: Although we used predicate definitions created by RE dataset creators, they are nevertheless a single symbolic and discrete form. We wondered about how the scores would change if we rephrased them by prompting GPT-4 to “rewrite” them without any special instructions. We also tested a simple baseline that just has the predicate name and entity spans fill the slots in this template: @predicate@: @subject@, @object@.
@@
Pipeline setup with an NER tool for end-to-end evaluation: Without assuming the availability of gold entity spans, we evaluate our model in an end-to-end manner using a pipeline that connects an NER tool’s output to our relation classifier. Specifically, we first apply the span-based PURE model 14 to identify entities, and then use these predictions as inputs to our relation extraction model (Figure 2). This reflects a realistic scenario enabling us to assess the model’s robustness under practical conditions. Since relation extraction depends on accurate entity boundaries and types, errors from the NER stage may propagate and affect overall performance. By adopting the PURE NER component, a strong baseline aligned with prior work, we ensure a fair evaluation. The comparison results in this pipeline setting are presented in Table 6. Our model, when combined with the PURE NER component, achieves state-of-the-art performance, surpassing previous approaches in the end-to-end setting. This demonstrates that the pipeline setup, despite relying on predicted entities, remains highly effective and outperforms existing joint models 25; 26.
@@
Next, consider this input from the ChemProt dataset: “Down-regulation of prostate-specific antigen (PSA) expression, an AR-target gene, by estramustine and bicalutamide was accompanied by the blockade of the mutated androgen receptor.” The model was able to identify the INDIRECT-DOWNREGULATOR relation between drugs estramustine and bicalutamide (subjects) and the protein PSA (object). But it failed to spot the same relation of those drugs with the object protein AR, which appears to be implicitly stated. Considering ChemProt extraction sometimes involves the full abstract, other sentences surrounding this may offer indirect clues about the relation. However, upon examining the full abstract, we are not able to see stronger evidence than what is already present in the sentence shown here.
Another common error pattern is due to the model’s lack of grasp of deep domain knowledge, particularly in biomedical datasets when information is densely packed. For example, consider this pithy ChemProt input: “In vivo, agonist actions of yohimbine at 5-HT(1A) sites are revealed by WAY 100,635-reversible induction of hypothermia in the rat.” Here the gold relation is the ANTAGONIST link between subject WAY 100,635 and object 5-HT(1A). The agonist actions of the chemical yohimbine on 5-HT(1A) result in hypothermia and the fact that this hypothermia can be reversed by WAY 100,635 indicates that it is playing an antagonist role for 5-HT(1A). This is a complex expression involving the intermediate entity yohimbine and an unusual looking phrase WAY 100,635-reversible that densely packs meaning that typically needs a new sentence to convey explicitly. This may simply be a case of a highly complex example needing multi-hop reasoning that needed to be carried out with a compact input text.